OpenStudy (anonymous):

Can anyone show me how this converges, please? (2n-1)!/(2n+1)!

OpenStudy (anonymous):

? I suppose, \[\sum^b_{n=0}\frac{(2n+1)!}{(2n+1)!}=\sum^b_{n=0}1=b\]

OpenStudy (anonymous):

Oh I see lol I made a sign mistake I give up I got to go to bed lol

OpenStudy (anonymous):

(2n-1)!/(2n+1)(2n)(2n-1)!

hartnn (hartnn):

so your function actually is 1/(2n+1)(2n) = 1/ (4n^2 +2n) use any test to show that this converges.

OpenStudy (anonymous):

k bro